A lab sample of gas is taken through cycle abca .The net work done on the gas is -1.2J.Along path ab,the change in internal energy is +3.0J and the work done on the gas is -5.0J.Along path ca,the energy transferred to the gas as heat is +2.5J.




(a). Determine the energy transferred to the gas as heat along path ab.



(b). Determine the energy transferred to the gas as heat along path bc.

1
Expert's answer
2021-11-24T19:16:39-0500



"∆W_{ca} = -1.2\textsf{ J}\\
∆U_{ab} = +3.0\textsf{ J}\\
∆W_{ab} = -5.0\textsf{ J}\\
Q_{ca} = +2.5\textsf{ J}\\"


According to the first law of thermodynamics;

"ΔU_{ca} = Q_{ca} − ∆W_{ca}\\
+2.5= Q_{ca}-(-1.2)\\
2.5 = Q_{ca}+1.2\\
Q_{ca} = 2.5-1.2\\
Q_{ca} = 1.3\textsf{ J}"


(a)

According to the first law of thermodynamics;

"Q_{ab} =\ ?\\
\quad\\
ΔU_{ab} = Q_{ab} − ∆W_{ab}\\
3.0 = Q_{ab} -(-5.0)\\
3.0 = Q_{ab} + 5.0\\
Q_{ab}= 3.0- 5.0 = -2.0\textsf{ J}\\"


(b)

Heat is a path function

"Q_{bc} =\ ?\\
\quad\\
1.3\textsf{ J}-2.0\textsf{ J} + Q_{bc} = 0\\
Q_{bc} = 0.7\textsf{ J}"